4. In January of 2000, the Dow Jones Index topped at 11,500. Using the January 2020 Dow of 29,000, what is the compound annual growth rate of an investment in this major index? Does this seem like an attractive rate of return? Does it indicate that the market is over or under valued?
Answer #1

In January of 2000, the Dow Jones Index topped at 11,500. Using the January 2020 Dow of 29,000, what is the compound annual growth rate of an investment in this major index? Does this seem like an attractive rate of return? Does it indicate that the market is over or under valued?

Sol. - In January of 2000 value of Dow Jones index = 11500

In January of 2020 value of Dow Jones index = 29000

Time = 20 year (2000-2020)

Future valve = present value * (1+InterestRate)time

29000 = 11500 * (1+i)20

(1+i)20 = 29000/11500 = 2.522

i = (2.522)1/20-1   = 1.0473-1   = .0473   or 4.73%

So, 4.73 is the compound annual growth rate of an investment in Dow Jones. Ans

This seem like an attractive rate of return because of currently United State interest rate is 1.75% fed. Treasury Rate &(4.6% approx. 20 year treasury rate mean) but Dow Jones give better return with respect to this. Ans.

Currently Dow Jones index rate of return of more then treasury rate that high in value but since 2000 to 2020 PE ratio of the index is decline from 26 to 24(estimated as per data ) that indicate Dow Jones is slightly as per par its not over valued or under valued. Ans.